CAVITY RING-DOWN SPECTROSCOPY OF THE 3-0 QUADRUPOLE OVERTONE BAND OF HYDRO-GEN

\underline{DANIEL C. ROBIE}, Chemistry Department, Barnard College, New York City, NY 10027; JOSEPH T. HODGES, J. PATRICK LOONEY, and ROGER D. VAN ZEE, Chemical Sciences and Technology Laboratory, National Institute of Standards and Technology, Gaithersburg, MD 20899.

We have used cavity ring-down spectroscopy to observe several lines in the 3-0 band of the quadrupole overtone spectrum of H_{2}. At pressures up to 2 bar in a cavity 1 m long, we have baseline losses less than 30 ppm and a signal-to-noise ratio above 40. New measurements of the line positions, linestrengths, pressure-broadening and pressure-shift coefficients will be presented, and compared with theoretical calculations.